two tears ago, Anton was three times as old as his sister, Nathalie. in 4 years, he will be only twice his sister's age. find their present ages.
Answer by mananth(16946)

two years ago, Anton was three times as old as his sister, Nathalie. in 4 years, he will be only twice his sister's age. find their present ages.
...
let Nathalie's age 2 years ago be x
Anton 2 years ago was 3x
...
In 4 years
Nathalie's age = x+6
Anton 's age = 3x+6
..
3x+6 = 2(x+6)
3x+6 = 2x+12
3x-2x = 12-6
x= 6 Nathalie's age 2 years ago. So now she will be 8 years
Anton was 3x years 2 years ago = 3*6 =18
so Now he is 18+2 = 20 years
..
m.ananth@hotmail.ca